OpenStudy (anonymous):

what is the equation of the line which includes points (0,3) and (4,5)? A. y= -1/2x+3 B. y=2x+3 C. y=-2x+3 D. y=1/2x+3

OpenStudy (anonymous):

plug x = 4 into each equation and see if you get y=5

OpenStudy (anonymous):

i got B for my answer

OpenStudy (anonymous):

is tht right?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

D is the right answer. y=1/2(4)+3 =2+3 =5